sydneymadden:

BJ's makes low carb cookies that cost $2.00 each. BJ's expects 15% of the cookies to fall apart and be discarded. BJ's wants a 45% markup on cost and produces 200 cookies. What should BJ's price each cookie? Round to the nearest cent.

Markthegoat1117:

To determine the price BJ's should set for each cookie, we need to consider the cost, the expected percentage of discarded cookies, and the desired markup. First, let's calculate the cost per cookie. BJ's makes low carb cookies that cost $2.00 each. Therefore, the cost per cookie is $2.00. Next, we need to calculate the number of cookies that will be discarded. BJ's expects 15% of the cookies to fall apart and be discarded. Since BJ's produces 200 cookies, the number of cookies expected to be discarded is 15% of 200, which is 0.15 * 200 = 30 cookies. Now, let's calculate the total cost of the cookies that will be sold. Since 30 cookies will be discarded, BJ's will sell 200 - 30 = 170 cookies. To determine the desired markup, BJ's wants a 45% markup on cost. The markup is calculated by multiplying the cost per cookie by the markup percentage: $2.00 * 0.45 = $0.90. Finally, we can calculate the price BJ's should set for each cookie. The price per cookie is the cost per cookie plus the markup: $2.00 + $0.90 = $2.90. Therefore, BJ's should price each cookie at $2.90, rounded to the nearest cent.

toga:

To determine the price of each cookie, we need to first calculate the cost of producing each cookie. Since BJ's produces 200 cookies and expects 15% to be discarded, they will only be able to sell 170 cookies (200 x 0.85 = 170). To calculate the cost per cookie, we need to divide the total cost of production by the number of cookies produced. The cost of producing 170 cookies can be calculated as follows: Cost of production = $2.00 x 170 = $340.00 To determine the markup price, we need to add the desired markup percentage to the cost per cookie: Markup price = Cost per cookie + (Cost per cookie x Markup percentage) Markup percentage = 45% Cost per cookie = $340.00 / 170 cookies = $2.00 Markup price = $2.00 + ($2.00 x 0.45) = $2.90 Therefore, BJ's should price each cookie at $2.90 to achieve a 45% markup on cost.